IDENTIFYING FLUORESCENT HYDROCARBON RADICALS FROM A BENZENE DISCHARGE

NEIL J. REILLY, DAMIAN L. KOKKIN, SCOTT H. KABLE and \underline{TIMOTHY W. SCHMIDT}, School of Chemistry, University of Sydney, NSW 2006, Australia.

We have applied the tools of laser induced fluorescence and dispersed fluorescence to the products of a benzene discharge in a free jet with a view to identifying new hydrocarbon radicals. We have observed several new band systems in the visible and UV regions, including a putative origin at 4759Å(air) which fluoresces to a ground state exhibiting vibrational modes at 120 cm^{−1}, 821 cm^{−1}, 988 cm^{−1} and 1160 cm^{−1}. This species is of particular interest due to its proximity to the centre of the (very) diffuse interstellar band at 4761.7Å (FWHM 25Å).